OpenStudy (anonymous):

Find the product. Simplify if possible. -1/3 * (8 1/7)

OpenStudy (anonymous):

a/b times c/d = ac/bd

OpenStudy (anonymous):

What about the 8? That's a whole number. What do I do?

OpenStudy (anonymous):

a times b/c = ab/c

OpenStudy (campbell_st):

create the improper fraction 8 1/7 = ( 8 x 7 + 1)/ 7 so the improper fraction is 57/7

OpenStudy (campbell_st):

so your problem is -1/3 * 57/7

OpenStudy (anonymous):

So the answer would be.. -2 5/7?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

oh, I see, it is \[8 + 1/7\]